OpenStudy (gabylovesyou):

The population of a type of local dragonfly can be found using an infinite geometric series where a1 = 65 and the common ratio is one sixth. Find the sum of this infinite series that will be the upper limit of this population. @phi

OpenStudy (gabylovesyou):

78 28 11 32

OpenStudy (phi):

do your notes have a formula for the sum of a geometric series?

OpenStudy (anonymous):

\[\left| r \right|=\frac{ 1 }{ 6 }<1\] \[s=\frac{ a }{ 1-r }\]

OpenStudy (anonymous):

a=65

OpenStudy (gabylovesyou):

@surjithayer i still dont know how to find r

OpenStudy (gabylovesyou):

@FibonacciChick666

OpenStudy (fibonaccichick666):

r is the common ratio

OpenStudy (gabylovesyou):

SO ... \[s = \frac{ 65 }{ 1 - 1/6 }\]

OpenStudy (gabylovesyou):

so the answer is 78 @FibonacciChick666 ?

OpenStudy (anonymous):

r is given as 1/6 your answer is correct.

OpenStudy (anonymous):

\[when ~\left| r \right|<1,r^n \rightarrow0 ~as~n \rightarrow \infty \]

OpenStudy (gabylovesyou):

thank you.

OpenStudy (anonymous):

yw